OpenStudy (anonymous):

Use the part of the quadratic formula that you chose above and find its value given the following quadratic equation: 2x2 + 7x + 3 = 0

jhonyy9 (jhonyy9):

- so this is a math question ,but because i have got it here i will add to you an answer here : 2x^2 +7x +3 =0 discriminant D=7^2 -4*2*3 =49-24 =25 x_1,2=(-7+/- sqrt25)/4 =(-7+/-5)/4 =-12/4 =-3 and -2/4 =-1/2 x_1=-3 x_2=-1/2 - hope so much that is understandably sure and was my pleasure all your help - good luck - bye

OpenStudy (trancenova):

Yup is a math question, please post questions in the right section :)